# Resilience for Regular Path Queries: Towards a Complexity Classification

### v1

@@ Abstract.
We show that computing the resilience in this sense is tractable (even in combined complexity) for all RPQs defined from so-called local languages. By contrast, we show hardness in data complexity for RPQs defined from the following language classes (after reducing the languages to eliminate redundant words): all finite languages featuring a word containing a repeated letter, and all languages featuring a specific kind of counterexample to being local (which we call four-legged languages). The latter include in particular all languages that are not star-free. Our results also imply hardness for all non-local languages with a so-called neutral letter.
We also highlight some remaining obstacles towards a full dichotomy. In particular, for the RPQ $abc|be$, resilience is tractable but the only PTIME algorithm that we know uses submodular function optimization.
@@ 1. Introduction
This paper studies the resilience problem which asks, for a query $Q$, a given database $D$ (in set or bag semantics), and a query answer $\mathbf{t}$, what is a minimum-cardinality set of tuples to remove from $D$ so that $\mathbf{t}$ is no longer a result to $Q$. (In particular: if $\mathbf{t}$ is not an answer of $Q$ on $D$ then resilience is zero; if $Q$ returns $\mathbf{t}$ on $D$ and all subinstances of $D$ then resilience is infinite.) The notion of resilience is motivated by the need to identify how “robust” an answer is, especially when the facts of the input database may be incorrect. Specifically, resilience quantifies how many tuples of $D$ must be removed before $\mathbf{t}$ is no longer a result, so that higher resilience values mean that a query result is less sensitive if tuples are removed.
The resilience problem has been introduced for conjunctive queries and mostly studied for such queries (DBLP:journals/pvldb/FreireGIM15; DBLP:conf/pods/FreireGIM20; bodirsky2024complexity; makhija2024unified). In this work, we investigate the computational complexity of resilience, and thus focus on the resilience of Boolean queries $Q$, i.e., how many tuples of the database $D$ must be removed to make $Q$ false. The key challenge in this line of work is to understand the data complexity of the problem, i.e., its complexity for a fixed query $Q$, as a function of the input database $D$. More precisely, the resilience problem can be tractably solved for some queries (e.g., trivial queries), and is known to be NP-hard for some other queries (see, e.g., (makhija2024unified)). Membership in NP always holds whenever query evaluation for $Q$ can be done in polynomial data complexity.
@@
In this paper, we study the resilience problem from a different perspective: we focus on graph databases (i.e., databases on an arity-two signature), and on queries which are defined by regular languages, namely, Boolean Regular Path Queries (Boolean RPQs), with no constants and with existentially quantified endpoints. In other words, the query asks whether the database contains a walk whose edge labels forms a word of the regular language, and resilience asks how many database facts must be removed so that no such walk exists. To our knowledge, we are the first to study this question, with one important exception: the recent paper of Bodirsky et al. (bodirsky2024complexity) which studies resilience via CSP techniques. Their results apply to much more general languages than CQs, and in particular imply a dichotomy on the resilience problem for RPQs, but in the specific setting of bag databases and with no insight on the combined complexity or on the criteria distinguishing tractable and intractable queries; we discuss this in more details at the end of Section 2.
@@
Our study of resilience for RPQs is motivated by database applications, but it is also motivated by connections to classical network flow problems. Consider for instance the classical MinCut problem of finding the minimum-cost cut in a flow network. The MinCut problem is easily seen to be equivalent to the resilience problem for the RPQ $ax^{*}b$ under bag semantics, with $a$-facts of the database corresponding to sources, $b$-facts to sinks, and $x$-facts to edges of the network. Hence, the resilience problem for RPQs corresponds to a generalization of MinCut, where the edges of the flow network are labeled and we must cut edges to ensure that no walk in a specific language remains. In the other direction, for many RPQs, we can devise tractable algorithms for resilience by reducing to a MinCut problem on a suitably-constructed flow network. However, resilience is NP-hard for some RPQs, e.g., the finite RPQ $aa$ (DBLP:conf/pods/FreireGIM20), so these methods do not always apply. Our results aim at determining what are the RPQs for which resilience can be tractably computed (in particular via such methods), and what are the RPQs for which it is intractable.
@@ 2. Preliminaries and Problem Statement
Observe that $Q_{L}$ does not change if we remove from $L$ all words $\beta$ such that some strict infix $\alpha$ of $\beta$ is also in $L$: indeed if $D$ contains a $\beta$-walk then it also contains an $\alpha$-walk. Formally, the reduced language of $L$ is $\mathtt{reduce}(L)\coloneq\{\alpha\in L\mid\text{no strict infix of }\alpha\text{ is in }L\}$. We then have that $Q_{L}$ and $Q_{\mathtt{reduce}(L)}$ are the exact same queries. The reduced language is also called the infix-free sublanguage of $L$ in (barcelo2019boundedness). We say that $L$ is reduced if $L=\mathtt{reduce}(L)$.
@@
We also consider the resilience problems on bag graph databases, by opposition with set graph databases which are the databases defined so far. A bag (graph) database consists of a database $D$ together with a multiplicity function $\mathtt{mult}:D\to\mathbb{N}_{>0}$; it can equivalently be seen as a multiset of facts. A Boolean query $Q$ is satisfied by a bag database $(D,\mathtt{mult})$ iff $D$ satisfies $Q$, i.e., the multiplicities of facts are not seen by queries. A contingency set for $Q$ and $D$ is then a subset $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$, and the resilience of $Q$ on $D$ is: $\mathtt{RES}_{\mathrm{bag}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}\sum_{e\in D^{\prime}}\mathtt{mult}(e)$, i.e., fact multiplicities represent costs.
@@
In this paper, we study the resilience problem for queries $Q_{L}$ defined by languages $L$; most of our results are for regular languages. The resilience problem for $L$, written $\mathtt{RES}_{\mathrm{set}}(L)$, is the following: given as input a database $D$ and $k\in\mathbb{N}$, decide if $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)\leq k$. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ in bag semantics is defined analogously. Note that $\mathtt{RES}_{\mathrm{set}}(L)$ easily reduces to $\mathtt{RES}_{\mathrm{bag}}(L)$ by giving multiplicity of 1 to each fact; however, in the setting of CQs, some conjunctive queries $Q_{0}$ are known for which $\mathtt{RES}_{\mathrm{set}}(Q_{0})$ is in PTIME, but $\mathtt{RES}_{\mathrm{bag}}(Q_{0})$ is NP-hard (makhija2024unified, Table 1).
@@
As discussed in the introduction, Theorem 2.2 gives a complete dichotomy across all regular languages, but it also has some shortcomings. (1.) First, the VCSP approach gives no explicit characterization of the class $L$ of tractable languages, or efficient bounds on the complexity of testing whether a given $L$ is PTIME or not: the criterion in (bodirsky2024complexity), using (kolmogorov2019testing), can be computationally checked but takes time exponential in a representation of the dual of $L$. (2.) Second, in tractable cases, the VCSP approach does not give good bounds on the degree of the polynomial (unlike, e.g., explicit reductions to MinCut). We remark that difficulties (1.) and (2.) also occur in works studying other problems where VCSP implies a general dichotomy but without explicit characterizations or tight bounds: this is the case for instance in (hirai2021minimum). (3.) Third, the VCSP approach does not shed light on the combined complexity, i.e., it does not identify cases that are tractable when both the query and database are given as input. (4.) Fourth, the VCSP approach inherently only applies to bag semantics and not to set semantics, so there is no known analogous dichotomy for $\mathtt{RES}_{\mathrm{set}}$.
@@ 3.1.2. Equivalence Between Local Languages and Letter-Cartesian Languages
If the input language $L$ is given as an NFA, we do not know the complexity of testing if $L$ is local: we show in Section 10 that a related problem is PSPACE-hard. Another question that remains open is to test whether the reduction of a given language $L$ is local: we can construct an automaton $A^{\prime}$ representing $\mathtt{reduce}(L)$ from an automaton $A$ representing $L$ and then study $A^{\prime}$, but the straightforward construction for $A^{\prime}$ entails an exponential blowup ((barcelo2019boundedness, Proposition 6)).
@@ 4.1. Hardness of $aa$
Let us first show that resilience is hard for the language $aa$. The result is known from the setting of CQs (DBLP:conf/pods/FreireGIM20), but we give a self-contained proof to sketch the constructions that come next.
@@
The resilience of $aa$ on $\Xi$ is then equal to the vertex cover number of a 5-subdivision [^4] A simpler construction can achieve a 3-subdivision, but our construction will generalize better for subsequent definitions. of $G$. Indeed, define the graph of $aa$-matches on a database $D$ as the undirected graph having one vertex for each fact of $D$ and one edge for each pair of facts of $D$ that forms an $aa$-path. It is then clear that $\mathtt{RES}_{\mathrm{set}}(aa,\Xi)$ is equal to the vertex cover number of the graph of $aa$-matches of $\Xi$. Moreover, the graph of $aa$-matches of each $D_{e}^{\prime}$ is a path of length 5 from vertex $F_{e,\textup{in}}$ to vertex $F_{e,\textup{out}}$ (see Figure 3(b)). Thus, the graph of $aa$-matches of $\Xi$ is a 5-subdivision of $G$, and we conclude by Section 4.2. ∎
@@ 4.4. Hardness Gadgets
To show the hardness of resilience for other languages than $aa$, we will build hardness gadgets, or gadgets for brevity. Gadgets are used as building blocks to reduce from the vertex cover problem, generalizing the idea of Figure 3(a) from Section 4.1. We first define pre-gadgets as databases with two domain elements to which endpoint facts will be attached. We then define the encoding $\Xi$ of a directed graph $G$ by gluing together copies of the pre-gadgets following $G$, like in the proof of Proposition 4.1. We last define gadgets as pre-gadgets whose hypergraph of matches of the query $Q$ forms an odd chain. The vertex cover problem on $G$ then reduces to the resilience of $Q$ on $\Xi$.
@@
For now, our definition of pre-gadgets is independent from the query. Let us now formalize what we require about query matches in gadgets, in order to reduce from the vertex cover problem. We do this by introducing the hypergraph of matches, generalizing the graph of matches from the proof of Proposition 4.1, along with a condensation operation to simplify it.
@@ 4.6. Condensed Hypergraph of Matches
Let $L$ be a language and $D$ a database. A match (or witness) of $L$ on $D$ is a set of edges corresponding to an $L$-walk (as defined in Section 2), i.e., a set $\{e_{1},\ldots,e_{m}\}$ such that $e_{1},\ldots,e_{m}$ is an $L$-walk. Then:
@@
Hence, from now on, all hardness results will be shown by simply exhibiting hardness gadgets for the languages in question. We also wrote an implementation (gadgetrepo) to check that gadgets are correct.
@@ 5.1. Four-Legged Languages
Let $L$ be a reduced language which is four-legged, and let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be legs that witness it, i.e., $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but $w^{\prime}\coloneq\alpha^{\prime}x\delta^{\prime}\notin L$.
If $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}$ are already stable (i.e., no infix of $w^{\prime}$ is in $L$) then we are done. Otherwise, as $w^{\prime}\notin L$, it means that there is a strict infix $w$ of $w^{\prime}$ that is in $L$. We know that $w$ must have a non-trivial intersection with $\alpha^{\prime}$, as otherwise $w\in L$ would be a strict infix of $\gamma^{\prime}x\delta^{\prime}\in L$, contradicting the fact that $L$ is reduced. Similarly, $w$ has a non-trivial intersection with $\delta^{\prime}$, as otherwise $w\in L$ would be a strict infix of $\alpha^{\prime}x\beta^{\prime}\in L$. So let us write $\alpha^{\prime}=\alpha_{2}\alpha_{1}$ and $\delta^{\prime}=\delta_{1}\delta_{2}$, with $\alpha_{1}$ and $\delta_{1}$ being non-empty, so that $w=\alpha_{1}x\delta_{1}$ is a strict infix of $\alpha^{\prime}x\delta^{\prime}$ that is in $L$. Note that $\alpha_{2}$ and $\delta_{2}$ cannot both be empty, as otherwise we have $w=w^{\prime}$ but we assumed that $w$ is a strict infix of $w^{\prime}$.
@@
For completeness, we point out that Section 5.2 does not hold if $L$ is not reduced. Indeed, take $L=\{x,axb,cxd\}$, which is four-legged, with the only possible body being $x$ and legs $(\alpha,\beta,\gamma,\delta)=(a,b,c,d)$; but these legs are not stable because $x$ is a infix of $\alpha x\delta=axd$ that is in the language.
@@
Note that Section 5.2 is not restricted to regular languages and applies to arbitrary languages (even when resilience is not even in NP). Next, we spell out two consequences of the result.

### v2

@@ Abstract.
We show that computing the resilience in this sense is tractable (even in combined complexity) for all RPQs defined from so-called local languages. By contrast, we show hardness in data complexity for RPQs defined from the following language classes (after reducing the languages to eliminate redundant words): all finite languages featuring a word containing a repeated letter, and all languages featuring a specific kind of counterexample to being local (which we call four-legged languages). The latter include in particular all languages that are not star-free. Our results also imply hardness for all non-local languages with a so-called neutral letter. We last highlight some remaining obstacles towards a full dichotomy.
@@ 1. Introduction
This paper studies the resilience problem which asks, for a Boolean query $Q$ and a given database $D$ (in set or bag semantics) what is a minimum-cardinality set of tuples to remove from $D$ so that $Q$ is no longer satisfied. (In particular: if $Q$ does not hold on $D$ then resilience is zero; if $Q$ holds on $D$ and all subinstances of $D$ then resilience is infinite.) The notion of resilience is motivated by the need to identify how “robust” a query is, especially when the facts of the input database may be incorrect. Specifically, resilience quantifies how many tuples of $D$ must be removed before $Q$ no longer holds, so that higher resilience values mean that $Q$ is less sensitive to tuples of $D$ being removed.
The resilience problem has been introduced for Boolean conjunctive queries and mostly studied for such queries (DBLP:journals/pvldb/FreireGIM15; DBLP:conf/pods/FreireGIM20; bodirsky2024complexity; makhija2024unified). The key challenge in this line of work is to understand the data complexity of the problem, i.e., its complexity for a fixed query $Q$, as a function of the input database $D$. More precisely, the resilience problem can be tractably solved for some queries (e.g., trivial queries), and is known to be NP-hard for some other queries (see, e.g., (makhija2024unified)). Membership in NP always holds whenever query evaluation for $Q$ can be done in polynomial data complexity.
@@
In this paper, we study the resilience problem from a different perspective: we focus on graph databases (i.e., databases on an arity-two signature), and on queries which are defined by regular languages, namely, Boolean Regular Path Queries (Boolean RPQs), with no constants and with existentially quantified endpoints. In other words, the query asks whether the database contains a walk whose edge labels forms a word of the regular language, and resilience asks how many database facts must be removed so that no such walk exists. To our knowledge, the only other work which studies this problem is the recent paper of Bodirsky et al. (bodirsky2024complexity), in which the authors use CSP techniques. Their results apply to much more general languages than CQs, and in particular imply a dichotomy on the resilience problem for Boolean RPQs (and also for the broader class of so-called two-way RPQs aka 2RPQs). However, their lower bounds do not apply to databases in set semantics. In addition, their dichotomy criterion, while decidable, does not give a intuitive language-theoretic characterization of the tractable cases. We discuss this in more details at the end of Section 2.
@@
Our study of resilience for RPQs is motivated by database applications, but it is also motivated by connections to classical network flow problems. Consider for instance the classical MinCut problem of finding the minimum-cost cut in a flow network. The MinCut problem, allowing multiple sources and sinks without loss of generality, is easily seen to be equivalent to the resilience problem in bag semantics for the RPQ $ax^{*}b$ on the alphabet $\{a,b,x\}$, with $a$-facts of the database corresponding to sources, $b$-facts to sinks, and $x$-facts to edges of the network. Hence, the resilience problem for RPQs corresponds to a generalization of MinCut, where the edges of the flow network are labeled and we must cut edges to ensure that no walk in a specific language remains. In the other direction, for many RPQs, we can devise tractable algorithms for resilience by reducing to a MinCut problem on a suitably-constructed flow network. However, resilience is NP-hard for some RPQs, e.g., the finite RPQ $aa$ (DBLP:conf/pods/FreireGIM20), so these methods do not always apply. Our results aim at determining what are the RPQs for which resilience can be tractably computed (in particular via such methods), and what are the RPQs for which it is intractable.
@@ 2. Preliminaries and Problem Statement
Observe that $Q_{L}$ does not change if we remove from $L$ all words $\beta$ such that some strict infix $\alpha$ of $\beta$ is also in $L$: indeed if $D$ contains a $\beta$-walk then it also contains an $\alpha$-walk. Formally, the reduced language of $L$ is $\mathtt{reduce}(L)\coloneq\{\alpha\in L\mid\text{no strict infix of }\alpha\text{ is in }L\}$. We then have that $Q_{L}$ and $Q_{\mathtt{reduce}(L)}$ are the exact same queries. The reduced language of $L$ is also called the infix-free sublanguage of $L$ in (barcelo2019boundedness): if $L$ is regular then so is $\mathtt{reduce}(L)$. We say that $L$ is reduced if $L=\mathtt{reduce}(L)$.
@@
We also consider the resilience problems on bag graph databases, by opposition with set graph databases which are the databases defined so far. A bag (graph) database consists of a database $D$ together with a multiplicity function $\mathtt{mult}:D\to\mathbb{N}_{>0}$; it can equivalently be seen as a multiset of facts. In the input, the multiplicities are expressed as integers written in binary. A Boolean query $Q$ is satisfied by a bag database $(D,\mathtt{mult})$ iff $D$ satisfies $Q$, i.e., the multiplicities of facts are not seen by queries. A contingency set for $Q$ and $D$ is then a subset $D^{\prime}\subseteq D$ such that $D\setminus D^{\prime}\not\models Q$, and the resilience of $Q$ on $D$ is: $\mathtt{RES}_{\mathrm{bag}}(Q,D)\coloneq\min_{\begin{subarray}{c}D^{\prime}\subseteq D\\ \text{s.t. }D\setminus D^{\prime}\not\models Q\end{subarray}}\sum_{e\in D^{\prime}}\mathtt{mult}(e)$, i.e., fact multiplicities represent costs.
@@
In this paper, we study the resilience problem for queries $Q_{L}$ defined by languages $L$; most of our results are for regular languages. The resilience problem for $L$, written $\mathtt{RES}_{\mathrm{set}}(L)$, is the following: given as input a database $D$ and $k\in\mathbb{N}$, decide if $\mathtt{RES}_{\mathrm{set}}(Q_{L},D)\leq k$. The problem $\mathtt{RES}_{\mathrm{bag}}(L)$ in bag semantics is defined analogously. Note that $\mathtt{RES}_{\mathrm{set}}(L)$ easily reduces to $\mathtt{RES}_{\mathrm{bag}}(L)$ by giving multiplicity of 1 to each fact; however, in the setting of CQs, some conjunctive queries $Q_{0}$ are known for which $\mathtt{RES}_{\mathrm{set}}(Q_{0})$ is in PTIME, but $\mathtt{RES}_{\mathrm{bag}}(Q_{0})$ is NP-hard (makhija2024unified, Table 1) (even when the multiplicities of the input bag database are written in unary).
@@
We point out that this result can be extended to so-called two-way RPQs, and to the case where some relations are declared exogenous (i.e., they cannot be part of a contingency set). Hence, as discussed in the introduction, Theorem 2.2 gives a complete dichotomy across all regular languages, but it also has some shortcomings. (1.) First, the VCSP approach gives no explicit characterization of the class $L$ of tractable languages, or efficient bounds on the complexity of testing whether a given $L$ is PTIME or not: the criterion in (bodirsky2024complexity), using (kolmogorov2019testing), can be computationally checked but takes time exponential in a representation of the dual of $L$. (2.) Second, in tractable cases, the VCSP approach does not give good bounds on the degree of the polynomial (unlike, e.g., explicit reductions to MinCut). We remark that difficulties (1.) and (2.) also occur in works studying other problems where VCSP implies a general dichotomy but without explicit characterizations or tight bounds: this is the case for instance in (hirai2021minimum). (3.) Third, the VCSP approach does not shed light on the combined complexity, i.e., it does not identify cases that are tractable when both the query and database are given as input. (4.) Fourth, the VCSP approach inherently only applies to bag semantics and not to set semantics, so there is no known analogous dichotomy for $\mathtt{RES}_{\mathrm{set}}$. Our results will improve on some of these shortcomings, though unlike (bodirsky2024complexity) they do not establish a full dichotomy.
@@ 3.1.2. Equivalence Between Local Languages and Letter-Cartesian Languages
If the input language $L$ is given as an NFA, we do not know the complexity of testing if $L$ is local: we show in Section 10 that it is PSPACE-hard to test if $L$ satisfies a property related to being local. Another question that remains open is to test whether the reduction of a given language $L$ is local: we can construct an automaton $A^{\prime}$ representing $\mathtt{reduce}(L)$ from an automaton $A$ representing $L$ and then study $A^{\prime}$, but the straightforward construction for $A^{\prime}$ entails an exponential blowup ((barcelo2019boundedness, Proposition 6)).
@@ 4.1. Hardness of $aa$
Let us first show that resilience is hard for the language $aa$. The result is known from the setting of CQs (DBLP:conf/pods/FreireGIM20, Prop. 10) by a direct reduction from 3-SAT. We give a self-contained proof where we reduce instead from minimum vertex cover: this proof serves as an intuitive presentation of the more general constructions that come next.
@@
The resilience of $aa$ on $\Xi$ is then equal to the vertex cover number of a 5-subdivision [^4] A simpler construction can achieve a 3-subdivision, but our construction will generalize better for subsequent definitions. of $G$. Indeed, define the graph of $aa$-matches on a database $D$ as the undirected graph having one vertex for each fact of $D$ and one edge for each pair of facts of $D$ that forms an $aa$-path. It is then clear that $\mathtt{RES}_{\mathrm{set}}(aa,\Xi)$ is equal to the vertex cover number of the graph of $aa$-matches of $\Xi$. Moreover, the graph of $aa$-matches of each $D_{e}^{\prime}$ is a path of length 5 from vertex $F_{e,\textup{in}}$ to vertex $F_{e,\textup{out}}$ (see Figure 3(b)). Thus, the graph of $aa$-matches of $\Xi$ is (isomorphic to) a 5-subdivision of $G$, and we conclude by Section 4.2. ∎
@@ 4.4. Hardness Gadgets
To show the hardness of resilience for other languages than $aa$, we will build hardness gadgets, or gadgets for brevity. Gadgets are used as building blocks to reduce from the vertex cover problem, generalizing the idea of Figure 3(a) from Section 4.1. We first define pre-gadgets as databases with two domain elements to which endpoint facts will be attached. We then define the encoding $\Xi$ of a directed graph $G$ by gluing together copies of the pre-gadgets following $G$, as in the proof of Proposition 4.1. We last define gadgets as pre-gadgets whose hypergraph of matches of the query $Q$ forms an odd chain. The vertex cover problem on $G$ then reduces to the resilience of $Q$ on $\Xi$.
@@
For now, our definition of pre-gadgets is independent of the query. Let us now formalize what we require from query matches in gadgets, in order to reduce from the vertex cover problem. We do this by introducing the hypergraph of matches, generalizing the graph of matches from the proof of Proposition 4.1, along with a condensation operation to simplify it.
@@ 4.6. Condensed Hypergraph of Matches
Let $L$ be a language and $D$ a database. A match (or witness) of $L$ on $D$ is a set of edges corresponding to an $L$-walk (as defined in Section 2): formally, each $L$-walk $e_{1},\ldots,e_{m}$ defines a match which is the set $\{e_{1},\ldots,e_{m}\}$, noting that several $L$-walks may define the same match. Then:
@@
Hence, from now on, all hardness results will be shown by simply exhibiting hardness gadgets for the languages in question. We also wrote an implementation (gadgetrepo) to automatically check some correctness properties on our gadgets.
@@ 5.1. Four-Legged Languages
Let $L$ be a reduced language which is four-legged, and let $x\in\Sigma$ be a body and $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}\in\Sigma^{+}$ be legs that witness it, i.e., $\alpha^{\prime}x\beta^{\prime}\in L$ and $\gamma^{\prime}x\delta^{\prime}\in L$ but $\eta^{\prime}\coloneq\alpha^{\prime}x\delta^{\prime}\notin L$.
If $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}$ are already stable (i.e., no infix of $\eta^{\prime}$ is in $L$) then we are done. Otherwise, as $\eta^{\prime}\notin L$, it means that there is a strict infix $\eta$ of $\eta^{\prime}$ that is in $L$. We know that $\eta$ must have a non-trivial intersection with $\alpha^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\gamma^{\prime}x\delta^{\prime}\in L$, contradicting the fact that $L$ is reduced. Similarly, $\eta$ has a non-trivial intersection with $\delta^{\prime}$, as otherwise $\eta\in L$ would be a strict infix of $\alpha^{\prime}x\beta^{\prime}\in L$. So let us write $\alpha^{\prime}=\alpha_{2}\alpha_{1}$ and $\delta^{\prime}=\delta_{1}\delta_{2}$, with $\alpha_{1}$ and $\delta_{1}$ being non-empty, so that $\eta=\alpha_{1}x\delta_{1}$ is a strict infix of $\alpha^{\prime}x\delta^{\prime}$ that is in $L$. Note that $\alpha_{2}$ and $\delta_{2}$ cannot both be empty, as otherwise we have $\eta=\eta^{\prime}$ but we assumed that $\eta$ is a strict infix of $\eta^{\prime}$.
@@
For completeness, we point out that Section 5.2 does not hold if $L$ is not reduced:
{remark}
Take $L=\{x,axb,cxd\}$, which is four-legged, with the only possible body being $x$ and legs $(\alpha,\beta,\gamma,\delta)=(a,b,c,d)$; but these legs are not stable because $x$ is a infix of $\alpha x\delta=axd$ that is in the language.
@@
Note that Section 5.2 is not restricted to regular languages and applies to arbitrary languages (even when resilience is not in NP). Next, we spell out two consequences of the result.